INAF LBC+LUCIFER queue observing, 2010 January 15-16
 
Observer: V. Testa, F. Mannucci, D. Thompson
Telescope Operator: WWack
Instrument: LUCIFER 
Data filename root: luci.20100116
 Summary 
* clear at the beginning of the night, thin and thicker clouds later; seeing  around 1.0-1.2" during the first half of he night, 0.5" later
* Lost Time: 15 min beginning of night for pointing; 30 min at 5:30 UT for lost collimation; 20 min at 10 UT for lost collimation 
  * 
Full logbook
  Data Summary 
	
		
			|  Files  | 
			 Project  | 
			 Object  | 
			 Notes  | 
		
	
	
		
			|  0186-0209  | 
			 Longhetti  | 
			 S2F1-142  | 
			 150K, 2h, seeing~0.9"  | 
		
		
			|  0216-0224  | 
			 Longhetti  | 
			 S2F1-142  | 
			 150K, 45min  | 
		
		
			|  0279-0283  | 
			 Turatto  | 
			 SN2009kn  | 
			 200HK, 25min, flexures  | 
		
		
			|  0342-0360  | 
			 Mannucci  | 
			 Q1422+22  | 
			 K, 29 min, probably useless  | 
		
		
			|  0284-0307  | 
			 Turatto  | 
			 SN2009kn  | 
			 200HK, 2.5h  | 
		
	
  Details 
 Darks
	
		
			 Files   | 
			 DITxNDIT  | 
		
	
	
		
			|  0002-0016  | 
			 2sec x 10  | 
		
		
			|  0017-0031  | 
			 10sec x 9  | 
		
		
			|  0032-0040  | 
			 15sec x 6  | 
		
		
			|  0042  | 
			 300sec x 1  | 
		
		
			|  0421-0425  | 
			 300sec x 1  | 
		
	
Sky flats
	
		
			|  Files  | 
			 Filter  | 
			 Counts  | 
		
	
	
		
			|  0043-0052  | 
			 Ks  | 
			 ~12000  | 
		
		
			|  0061-0071  | 
			 Ks  | 
			 ~7000  | 
		
		
			|  0072-0091  | 
			 H  | 
			 15000-7000  | 
		
		
			|  0092-0111  | 
			 J  | 
			 11000-5000  | 
		
		
			|  0113-0133  | 
			 Ks  | 
			 300  | 
		
		
			|  0134-0153  | 
			 H  | 
			 900  | 
		
		
			|  0154-0173  | 
			 J  | 
			 900  | 
		
	
01:00 - sky almost clear, a few thin clouds 
01:10 - Pointing, collimation 
01:40 - 
S2F1-142 - Longhetti, pointing 0180-0184 
01:55 - 
S2F1-142 : 150_K, 300sec X 24  0186-0210, seeing~0.9" 
04:18 - 
S2F1-142: check alignment (0210-0215) 
04:22 - 
S2F1-142 : 150_K, 1h (0216-0227): clouds towards the end of the integration, abort 
05:17 - Flat, halo1+halo2, dit=3, nexp=5 0225-0244 
05:22 - Arc Ar + Ne: dit=30 (0245-0248) 
05:26 - collimation lost while going to the std, possibly because of temperature jump of 4 deg. Pointing and collimation 
05:58 - Telluric STD: Hip 13917: grating=150K  (0257-0258). No collimation 
06:08 - Telluric STD: Hip 13917: grating=200H_K (0259-0260) 
06:27 - 
SN2009kn - Turatto pointing 0270-0278 
06:54 - 
SN2009kn : grating=200H+K, dit=300, 25min : variation in line position 
07:25 - 
SN2009kn: start again setting central_W=1.93 0284-0307: thin clouds at the end
09:45 - Flat, haol1+halo2,   0317-0326 
09:48 - Arc Ar + Ne: 0327-0330 
0956 - Telluric STD Hip37787: grating 200_H+K 0336-0337 
10:05 - problems with active control, elongated guide star, lost 20 min 
10:29 - 
Q1422 - Mannucci K, 0342-0360, seeing~0.5": aborted due to clouds (~1 mag ) 
11:25 - 
SN2009dc - Turatto, J, 0361-0420 
01:50 - Darks 0421-0425 
-- 
FilippoMannucci - 15 Jan 2010